Hello!
If a manufacturer sells x units of a product, his R=20x and C=2000 + 8x +.0025X^2 determine how many units he should sell to have maximum profit...
So far I know that P= R-C
So P = -.0025x^2 + 12x -2000
and
-b/2a=x soo... x=2400
how do i do the next step? I feel so dumb...
THanks so much

Answer by stanbon(75887)

Hello!
If a manufacturer sells x units of a product, his R=20x and C=2000 + 8x +.0025X^2 determine how many units he should sell to have maximum profit...
So far I know that P= R-C
So P = -.0025x^2 + 12x -2000
and
-b/2a=x soo... x=2400
----------------
Good job. And 2400 is the number of units
to sell to maximimize profit.
------------------
Cheers,
Stan H.